Consider the production function Q = 6L^2 - L^3.


4. Consider the production function Q = 6L^2 - L^3.

a. Find average product. Find the number of labor (L) that the firm should hire to maximize average product.

b. Find marginal product. Find the number of labor (L) that the firm should hire to maximize marginal product.

c. Find the number of labor (L) that the firm should hire to maximize total product of labor.

d. Find the number of labor (L) that the firm has zero average product.
Answer #1

Given Q= 6L2 - L3

a) APL= Q/L = 6L - L2

d.APl/dL = 6 - 2L

We put dAPL/dL = 0

So 6 - 2L = 0

So L = 3

b) MPL= 12L - 3L2

dMPl/dL = 12 - 6L

dMPl/dL = 0, 12-6L = 0

L = 2

c) TP= 6L2 - L3

dTP/dL = 12L - 3L2

dTP/dL= 0

So 12L - 3L2 = 0

So L= 4

d) AP = 6L - L2

AP = 0

6L - L2 = 0

L= 6